# PartGS: Learning Part-aware 3D Representations by Fusing 2D Gaussians and Superquadrics

### v2

@@ 3.2.1 Block-level Optimization
Rendering Loss. Based on the rendering process of Gaussian spaltting [18], the rendering loss is calculated by associating $L_{1}$ with a D-SSIM term[18] :
@@
Coverage Loss. This loss ensures that the block set $\{\mathcal{B}_{1}\dots\mathcal{B}_{M}\}$ covers the object while preventing it from extending beyond the object’s boundaries. To determine the 3D occupancy field based on the blocks, we first define the approximate signed distance of a 3D point $x$ to the $i$-th block, $D_{i}(x)=1-\Psi_{i}(x)$. Here, $\Psi_{i}$ denotes the superquadric inside-outside function [2] associated with the block $i$. Consequently, $D_{i}(x)\geq 0$ if the point $x$ lies inside the $i$-th block, and $D_{i}(x)<0$ if $x$ is outside the block. Inspired by NeRF [25], we sample a ray set $\mathcal{R}$ using ray-casting based on camera poses. Given the object mask, we associate each ray $r\in\mathcal{R}$ with a binary label: $l_{r}=1$ if the ray $r$ is inside mask, otherwise $l_{r}=0$. The coverage loss is defined as:
@@
Here, $L_{\text{cross}}$ encourages rays to intersect with blocks, while $L_{\text{non-cross}}$ discourages rays from intersecting with blocks. They are defined as:
$L_{\text{cross}}(r)=\text{ReLU}\Big{(}\min_{i\in\mathcal{M}}\min_{x_{j}^{r}\in{\mathcal{X}}_{r}}D_{j}(x_{j}^{r})\Big{)},$ (7)
@@
where $x_{j}^{r}$ means the $j$-th sampled points along the ray $r$ and =$\mathcal{M}=\{1,\dots M\}$. Intuitively, this means that there is at least one in sampled points ${\mathcal{X}}_{r}$ along the ray $r$ inside the mask that lies within a certain block, while all points along the ray outside the mask should not lie within any block.
Overlap Loss. We add a regularization term to encourage individual blocks to remain non-overlapping. Given the difficulty of directly calculating block overlap, we use a Monte Carlo method similar to  [27]. Practically, we sample multiple 3D points in space and penalize each point that lies inside more than $k$ blocks. Based on the definition of the superquadric inside-outside function, we devise a soft occupancy function:
$\mathcal{O}_{i}^{x}=\tau_{i}\text{ sigmoid}(\frac{1-\Psi_{i}(x)}{\gamma}),$ (9)
where $\gamma$ is a temperature parameter. Therefore, given a set of $N$ sampled 3D points $\Omega$, the overlap loss is expressed as:
@@
Parsimony Loss. To promote the use of the minimal number of blocks and achieve parsimony in decomposition, we add a loss term that regularizes the block opacity ($\tau$). This loss is defined as:
@@
Opacity Entropy Loss. Naturally, during optimization, the opacity of the block inside the object region tends to approach 1, while the opacity of the block outside the object region tends to approach 0. To facilitate this, we associate the opacities of the blocks with the labeled rays. Specifically, we use a cross-entropy loss $L_{ce}$, between the opacity and mask labels as follows:
@@
The total loss is the weighted summation of the loss terms described above:
@@
Adaptive Number of Blocks. Given that the number of parts varies across scenes, we allow the number of blocks to adjust during optimization adaptively. When the opacity of an existing block falls below a threshold $t$, this block is immediately removed. Additionally, we explore a block-adding mechanism using point clouds that are randomly generated within the scene area. Specifically, we use DBSCAN [34] to cluster initial point clouds (e.g., accessible by-products of COLMAP [33]) that are not covered by any blocks. A new block is then introduced at the center of each point cloud cluster. The other parameters of these new blocks are initialized randomly.
At last, block-level reconstruction gets a 3D scene composited by several superquadrics, as shown in Fig. 2.
3.2.2 Point-level Optimization
The primitive-based hybrid representation effectively decomposes the shape into parts but performs suboptimally for complex objects. To address this, we introduce a refinement stage for more precise geometry fitting. In this stage, we decouple the constraint between Gaussians and superquadrics, enabling the Gaussians to be freely optimized. Furthermore, to prevent Gaussian components from one block passing through another and disrupting the continuity and plausibility of the part decomposition, we impose a new constraint to minimize the negative signed distance of each Gaussian entering other blocks:
@@ 3.2.2 Point-level Optimization
where $\Omega$ is the sampled Gaussian set, and $\delta$ represents the identifier of the block to which the Gaussian $x$ belongs. As seen in Fig. 2, the reconstruction becomes more accurate and aligns better with the target shape after this stage.
@@
Figure 4: Qualitative results on BlenderMVS [45] and self-capatured data. We present RGB renderings and decomposed parts from novel views. The top examples are from the BlenderMVS dataset, and the last example is from our captured scenes.
@@
Ours (Point-level) Image 0.95 1.38 0.55 0.71 1.02 1.39 1.16 1.33 1.49 0.80 0.95 1.31 0.69 1.20 0.82 1.05 5.9
@@
Ours (Point-level) 0.95 35.04 0.99 0.015 $\sim$ 40 m
@@ 4 Experiments
We evaluate our approach on 3D reconstruction quality, view synthesis, shape parsimony, and reconstruction time, comparing it with previous state-of-the-art methods.
@@ 4.1 Datasets
We conduct comparative experiments on two public datasets: DTU [17] and ShapeNet [6]. To validate the effectiveness of our method in man-made objects, we construct a ShapeNet subset with four categories: Chair, Table, Gun, and Airplane. The DTU dataset comprises 15 commonly used scenes, and each category in the ShapeNet dataset contains 15 objects. Additionally, to demonstrate our method’s potential with real-world data, more qualitative results are presented on BlenderMVS dataset [45] and self-captured scenes.
@@ 4.2 Implementation Details
We build upon the 2DGS [16] architecture and add a non-learnable part attribute to each Gaussian component to enforce the constraint in Eq. 14. In the ShapeNet benchmark, we observe a common degradation in mesh reconstruction as shown in Fig. 7. To address this, we propose Gaussian scale regularization, as validated in Sec. 4.4. The same hyperparameters are used for all experiments. We set the initial number of primitives $M$ to 8. The temperature parameter $\gamma$ in Eq. 9 is 0.005 and the overlapping number $k$ in Eq. 10 is 1.95. We train the hybrid representation for 30k iterations, followed by refinement optimization for another 30k iterations. During block-level optimization, the block-adding operation is carried out at the 5k-th and 10k-th iterations. The loss weights $\lambda_{\text{cov}}$, $\lambda_{\text{over}}$, $\lambda_{\text{par}}$, and $\lambda_{\text{opa}}$ are set to 10, 1, 0.01, and 0.01, respectively. For more details on the dataset, metrics, baselines, and method, please refer to the supplementary materials.
@@ 4.3.1 DTU and Shapenet Benchmark
Geometry Reconstruction. In Tab. 1 and Tab. 3, we compare our geometry reconstruction to SOTA shape decomposition methods on Chamfer distance and training time using DTU and Shapenet dataset. Our method consistently outperforms in all scenes compared to prior works. As shown in Fig. 3, our approach consistently produces interpretable 3D decompositions with further refinement achieving more detailed geometry (the last two columns). MBF [31] achieves a low CD error but at the cost of using significantly more primitives, leading to over-decomposition and the inclusion of meaningless parts. Moreover, as shown in Tab. 2 and Tab. 4, our approach achieves competitive results with advanced Gaussian-based methods(e.g., 2DGS [16]) and surface reconstruction approaches (e.g., Neuralangelo [21]), which are limited to producing unstructured meshes. Notably, our model demonstrates excellent efficiency, with a reconstruction speed approximately 10X faster than PartNeRF and over 3X faster than DBW.
Appearance Reconstruction. Beyond part decomposition, our method enables high-fidelity image synthesis, attributed to integrating Gaussian splatting within superquadric representations. EMS and MBF lack image rendering capabilities, while PartNeRF and DBW yield low-quality view synthesis. In contrast, our method achieves high-quality appearance rendering results, as demonstrated in Tab. 2 and Tab. 4.
Figure 6: Qualitative comparison with SAM-based methods. Our method achieves finer-grained and more distinct part decomposition.
@@ 4.3.2 Real-life Data
To further demonstrate the applicability of our method for learning shape decomposition, we test our model on the real-life images from BlenderMVS dataset [45] and a self-captured dataset. As shown in Fig. 4, our approach can robustly produce both realistic appearances and meaningful 3D decompositions across a variety of data types. More results are provided in the supplementary material.
@@
Figure 9: Applications. Part-Aware Editing (top): After optimization, we can easily edit the scene by adding, scaling, or moving specific parts. 3D Content Generation (middle): By combining parts of different objects, we can create new 3D content. Physics simulations (bottom): The representation enables us to perform physics-based simulations easily.
@@ 4.3.3 Compared to SAM-based Methods
We also conduct comparisons to SAM-based methods for this task, as shown in Fig. 6. Despite impressive advances in 3D segmentation and editing [5, 19, 46, 47] by distilling 2D information, achieving finer-grained semantic disentanglement remains challenging due to indistinct textures and severe cross-view 2D inconsistencies. Lifting 2D segmentation features to 3D segmentation will also lead to feature misalignment, leading to visual artifacts. In contrast, our method segments directly from 3D space, enabling more fine-grained and clearly defined component decomposition.
@@ 4.4 Ablations
In this section, we first analyze the design choices by isolating each strategy to assess its impact. We report the averaged quantitative performance on 15 instances of chair category in Tab. 5 and provide visual comparisons in Fig. 5. Removing overlapping loss enhances reconstruction quality but reduces semantic clarity. Omitting opacity loss drastically affects the number of primitives and doubles the CD, highlighting its role in controlling primitive presence and reconstruction accuracy. Coverage loss is crucial in reconstruction accuracy, ensuring primitives align correctly with the target. The adaptive primitive strategy fills in missing parts and improves the reconstruction metric. Without parsimony loss, CD and rendering metrics are optimal but result in redundancy and over-decomposition.
We also validate the effectiveness of Gaussian scale regularization in Fig. 7. 2DGS [16] tends to use large-scale Gaussians in texture-less areas and create holes in the constructed mesh by TSDF. Holes occur since large-scale Gaussians can’t support consistent depth rendering from different views. Scale regularization effectively addresses this issue by suppressing large-scale Gaussians.
@@ 4.5 Applications
Fig. 9 illustrates three applications of our method that original 3DGS and NeRF-based approaches do not support. Firstly, after optimization, we obtain the part decomposition, facilitating easy editing of specific object or scene components, e.g., adding, moving, removing, or scaling. Secondly, by combining parts of different objects, our method enables the creation of new high-quality 3D content. Additionally, our method enables easy manipulation of objects within a scene, supporting physics-based simulations.
5 Conclusion
@@
We introduce a hybrid representation of superquadrics and 2D Gaussians, to learn 3D scenes in a part-aware representation. Compared to prior works, the proposed method also keeps geometry details instead of shape abstractions, supporting high-level image rendering quality. The method gets state-of-the-art performance in comprehensive evaluations. One limitation is that the method takes background-free scenes as inputs, from segmentation tools such as SAM, as we care about the structure of the centered scene the most. In the future, we aim to explore how to model backgrounds and extend the work to address entire complex scenes.

### v4

@@ 3.2.1 Optimization
Rendering Loss. The rendering loss is based on 3DGS [25], where it combines an $L_{1}$ term with a D-SSIM term:
@@
Coverage Loss. The coverage loss ensures that the block set $\{\mathcal{B}_{1}\dots\mathcal{B}_{M}\}$ covers the object while preventing it from extending beyond its boundaries. To determine the 3D occupancy field based on the blocks, we first define the approximate signed distance of a 3D point $x$ to the $i$-th block, $D_{i}(x)=\Psi_{i}(x)-1$. Here, $\Psi_{i}$ denotes the superquadric inside-outside function [2] associated with the block $i$. Consequently, $D_{i}(x)\leq 0$ if the point $x$ lies inside the $i$-th block, and $D_{i}(x)>0$ if $x$ is outside the block. Inspired by NeRF [34], we sample a ray set $\mathcal{R}$ using ray-casting based on camera poses. Given the object mask, we associate each ray $r\in\mathcal{R}$ with a binary label: $l_{r}=1$ if the ray $r$ is inside mask, otherwise $l_{r}=0$. The coverage loss is defined as:
@@
Here, $L_{\text{cross}}$ encourages rays to intersect with blocks, while $L_{\text{non-cross}}$ discourages rays from intersecting with blocks:
$L_{\text{cross}}(r)=\text{ReLU}\Big{(}\min_{i\in\mathcal{M}}\min_{x_{j}^{r}\in{\mathcal{X}}_{r}}D_{i}(x_{j}^{r})\Big{)},$ (6)
@@
where $x_{j}^{r}$ means the $j$-th sampled point along the ray $r$ and =$\mathcal{M}=\{1,\dots M\}$. Intuitively, this implies that at least one sampled points ${\mathcal{X}}_{r}$ along the ray $r$ inside the mask that lies within a certain block, while all points along the ray outside the mask should not belong to any block.
Overlap Loss. We introduce a regularization term to prevent overlap between individual blocks. Given the difficulty of directly calculating block overlap, we adopt a Monte Carlo method similar to  [36]. Specifically, we sample multiple 3D points in space and penalize those that lie inside more than $k$ blocks. Based on the superquadric inside-outside function, a soft occupancy function is defined as:
$\mathcal{O}_{i}^{x}=\tau_{i}\text{ sigmoid}(\frac{-D_{i}(x)}{\gamma}),$ (8)
where $\gamma$ is a temperature parameter. Thus, for a set of $N$ sampled 3D points $\Omega$, the overlap loss is expressed as:
@@
Parsimony Loss. To promote the use of the minimal number of blocks and achieve parsimony in decomposition, we introduce a regularization term that penalizes block opacity ($\tau$). This loss is defined as:
@@
Opacity Entropy Loss. During optimization, the opacity of the block inside the object region tends to approach 1, while the opacity of the block outside the object region tends to approach 0. To facilitate this, we associate the opacities of blocks with the labeled rays and apply a cross-entropy loss $L_{ce}$ between the block opacity and the mask labels, defined as follows:
@@
The total loss is the weighted sum of the loss terms described above:
@@
3.2.2 Adaptive Number of Blocks.
Given that the number of parts may vary across different scenes, we allow the number of blocks to adjust during optimization adaptively. Specifically, when the opacity of an existing block falls below a threshold $t$, this block is removed immediately. Furthermore, we explore a block-adding mechanism to dynamically incorporate new components, ensuring comprehensive coverage of target objects. Specifically, DBSCAN [45] is employed to cluster initial point clouds that are not covered by any blocks, where the point clouds are either randomly initialized or derived as by-products from COLMAP [44]. For each point cloud cluster, a new block is introduced at its center, with the remaining parameters of the block initialized randomly.
The final block-level reconstruction generates a textured 3D scene represented by multiple superquadrics, with surface details rendered through Gaussians, as illustrated in Fig. 2.
3.3 Point-level Decomposition
The primitive-based hybrid representation effectively decomposes the shape into parts but exhibits suboptimally for complex objects. To address this limitation, we perform a refinement stage that enhances geometric fitting. In this stage, the constraint between Gaussians and superquadrics is decoupled, enabling independent Gaussian optimization. Furthermore, to prevent Gaussian components from one block passing through adjacent blocks and disrupting the continuity and plausibility of the part decomposition, we impose a new constraint to minimize the negative signed distance of each Gaussian entering other blocks:
@@ 3.3 Point-level Decomposition
where $\Omega$ is the sampled Gaussian set, and $\delta$ represents the identifier of the block to which the Gaussian $x$ belongs. As seen in Fig. 2, the reconstruction becomes more accurate and aligns better with the target shape.
@@
Figure 4: Qualitative results on BlendedMVS [63] and self-captured data. We present RGB renderings and decomposed parts from novel views. The top examples are from the BlendedMVS dataset, and the last example is from our captured scenes.
@@
Ours (Point-level) Image 0.70 1.17 0.55 0.65 1.06 1.23 1.10 1.36 1.37 0.78 0.92 1.41 0.69 1.05 0.71 0.98 5.9
@@
Ours (Point-level) 0.98 35.04 0.99 0.015 $\sim$ 40 m
@@ 4 Experiments
We comprehensively evaluate our approach across four key aspects: 3D reconstruction quality, view synthesis performance, shape parsimony, and computational efficiency, comparing it with state-of-the-art methods in the field.
@@ 4.1 Datasets
Extensive experiments are conducted on two widely-used public datasets: DTU [23] and ShapeNet [6]. Specifically, 15 standard scenes from the DTU dataset that are widely adopted for reconstruction quality assessment are used for evaluation. Each scene contains either 49 or 64 images. Additionally, we construct a ShapeNet subset consisting of four categories: Chair, Table, Gun, and Airplane, with 15 objects selected per category. For each object, 50 rendered images are generated for training. We further validate real-world applicability on both the BlendedMVS dataset [63] and self-captured scenes, demonstrating robust performance across synthetic and natural environments.
@@ 4.2 Implementation Details
We build upon the 2DGS [21] architecture and add a non-learnable part attribute to each Gaussian component to enforce the constraint in Eq. 13. In the ShapeNet benchmark, we observe a common degradation in mesh reconstruction as shown in Fig. 7. To address this, we propose Gaussian scale regularization, as validated in Sec. 4.4. We set the initial number of primitives $M$ to 8. The temperature parameter $\gamma$ in Eq. 8 is 0.005 and the overlapping number $k$ in Eq. 9 is 1.95. We train the hybrid representation for 30k iterations, followed by refinement optimization for another 30k iterations. During block-level optimization, the block-adding operation is carried out at the 5k-th and 10k-th iterations. The loss weights $\lambda_{\text{cov}}$, $\lambda_{\text{over}}$, $\lambda_{\text{par}}$, and $\lambda_{\text{opa}}$ are set to 10, 1, 0.002, and 0.01, respectively. For more details about datasets, metrics, and baselines, please refer to Supplementary.
@@ 4.3.1 DTU and Shapenet Benchmark
Geometry Reconstruction. In Tab. 1 and Tab. 3, we compare our geometry reconstruction to SOTA shape decomposition methods on Chamfer distance and training time using DTU and Shapenet dataset. The Chamfer distance metrics reported for ShapeNet are scaled by a factor of 100 for readability. Our method consistently outperforms in all scenes compared to prior works. As shown in Fig. 3, our approach consistently produces interpretable 3D decompositions with further refinement achieving more detailed geometry (the last two columns). MBF [42] achieves a low CD error but at the cost of using significantly more primitives, leading to over-decomposition and the inclusion of meaningless parts. Moreover, as shown in Tab. 2 and Tab. 4, our approach achieves competitive results with the advanced Gaussian-based 2DGS [21] and surface reconstruction Neuralangelo [30], which are limited to producing unstructured meshes. Notably, our model demonstrates excellent efficiency, with a reconstruction speed approximately 10X faster than PartNeRF [48] and over 3X faster than DBW [36].
Appearance Reconstruction. In addition to part decomposition, our method enables high-fidelity image synthesis, attributed to integrating Gaussian splatting within hybrid representations. EMS [32] and MBF [42] operate directly on point clouds and consequently lack image rendering capabilities, while PartNeRF [48] and DBW [36] yield low-quality view synthesis. In contrast, our method achieves high-quality appearance rendering results, as demonstrated in Tab. 2 and Tab. 4.
Figure 6: Qualitative comparison with SAM-based methods. Visual results demonstrate that our method produces more structurally coherent decompositions, whereas SAM-based approaches frequently exhibit spatial discontinuities.
@@ 4.3.2 Real-life Data
To further demonstrate the applicability of our method for learning shape decomposition, we test our model on the real-life images from the BlendedMVS dataset [63] and a self-captured dataset. As shown in Fig. 12, our approach can robustly produce both realistic appearances and reasonable 3D decompositions across a variety of data types. More results are provided in the supplementary material.
@@
Figure 9: Applications. Part-Aware Editing (top): After optimization, we can easily edit the scene by adding, scaling, or moving specific parts. 3D Content Generation (bottom): By combining parts of different objects, we can create new 3D content.
@@ 4.3.3 Compared to SAM-based Methods
We also conduct comparisons to SAM-based [27] methods for this task, as shown in Fig. 6. Despite impressive advances in 3D segmentation and editing [5, 26, 65, 66] by distilling 2D information, achieving semantic disentanglement but remains challenging due to indistinct textures and severe cross-view 2D inconsistencies. Lifting 2D segmentation features to 3D segmentation will also lead to feature misalignment, leading to visual artifacts. In contrast, our method performs direct 3D space segmentation, yielding more reasonable and structurally coherent decompositions.
@@ 4.4 Ablations
In this section, we first analyze the design choices by isolating each strategy to assess its impact. We report the averaged quantitative performance on 15 instances of the chair category in Tab. 5 and provide visual comparisons in Fig. 5. Removing overlapping loss provides better reconstruction accuracy but more overlapping parts. The opacity loss drastically affects the number of primitives and doubles the CD, highlighting its role in controlling primitive presence and reconstruction quality. Coverage loss is crucial in reconstruction accuracy, ensuring primitives align correctly with the target. The adaptive primitive strategy fills in missing parts and improves the reconstruction metric. Without parsimony loss, CD and rendering metrics are optimal but result in over-decomposition. Note that the full model did not yield the highest accuracy because the part-aware reconstruction trades off the accuracy and rationality of the part decomposition. The objective is to achieve high accuracy while preserving reasonable parts.
We also validate the effectiveness of Gaussian scale regularization in Fig. 7. 2DGS [21] tends to use large-scale Gaussians in texture-less areas and create holes in the constructed mesh by TSDF integration [71]. The presence of holes stems from the inability of large-scale Gaussians to maintain view-consistent depth. Scale regularization effectively addresses this issue by suppressing large-scale Gaussians.
@@ 4.5 Applications
Fig. 9 illustrates two applications of our method that the original 3DGS and NeRF-based approaches do not support. Firstly, after optimization, we obtain the part decomposition, facilitating easy editing of specific objects or scene components, e.g., adding, moving, removing, or scaling. Secondly, by combining parts of different objects, our method enables the creation of new high-quality 3D content.
5 Conclusions
@@
We introduce PartGS, a hybrid representation of superquadrics and 2D Gaussians, to learn 3D scenes in a part-aware representation. Compared to prior works, the proposed method retains geometry details, supporting high-quality image rendering. It obtains state-of-the-art performance in comprehensive evaluations. One limitation is that it takes background-free scene images as inputs, leveraging segmentation tools such as SAM [27]. In the future, we aim to explore how to model backgrounds and extend to larger and more complex scenes.
Acknowledgements
This work is supported in part by the NSFC (62325211, 62132021, 62372457), the Major Program of Xiangjiang Laboratory (23XJ01009), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the Natural Science Foundation of Hunan Province of China (2022RC1104).
Supplementary Material
At first, this supplementary material provides detailed experimental settings, including data processing procedures, implementation details for comparison baseline methods, and evaluation metrics (Sec.6). Subsequently, we provide further implementation specifics, including the training configurations and rendering process (Sec.8).
In Fig. 10 and Fig. 11, we provide additional visual comparisons of our method against state-of-the-art baselines on DTU [23] and ShapeNet [6] datasets. Moreover, we include visual results on several scenes from the BlendedMVS [63] and self-captured data in Fig. 12. It can be observed that the decomposed reconstructions by our method are more reasonable and precise, which is believed to be beneficial for downstream tasks. Our code and data are available at https://github.com/zhirui-gao/PartGS.
6 More Details on Experiments
6.1 Datasets
We conduct evaluations on two public datasets: DTU [23] and ShapeNet [6]. DTU is a multi-view stereo (MVS) dataset comprising 80 forward-facing scenes, each captured with 49 to 64 images. In experiments, we use 15 publicly recognized scenes commonly adopted in previous studies [21, 30]. We downsample the images to a resolution of $400\times 300$ for computational efficiency.
To validate the effectiveness of our method in the decomposition of man-made objects, we construct a subset of the ShapeNet dataset comprising four categories: Chair, Table, Gun, and Airplane. Each category includes 15 distinct objects, providing diverse instances. For each object, we randomly sample 100 camera poses on a sphere and render images at a resolution of $400\times 400$. The rendered images are equally split into training and testing sets.
Additionally, to explore the potential of PartGS in handling real-life data, we present qualitative results on the BlendedMVS dataset [63] and self-captured scenes. For BlendedMVS, we use official camera poses, while for self-captured scenes, camera poses are estimated by COLMAP [44] and normalized using IDR [64]. All images are resized to $400\times 300$ pixels. For the DTU, ShapeNet, and BlendedMVS datasets, we utilize the ground-truth foreground masks provided within the datasets. For self-captured real-world scenes, we employ the Segment Anything Model (SAM) [27] to segment foreground objects.
6.2 Implementation Details on Baselines
We compare our method with four state-of-the-art works on 3D shape decomposition: EMS [32], MonteboxFinder (MBF) [42], PartNeRF [48] and DBW [36]. MBF and EMS are applied to point clouds, utilizing cuboids and superquadrics, respectively, to fit the 3D points. The input point clouds are sampled from either ground truth 3D shapes (GT) or meshes reconstructed by state-of-the-art MVS method Neus [55]. We adhere to the testing procedure outlined by DBW [36] and randomly sample 5K and 200K points from the GT meshes as the input for EMS and MBF, respectively. Similar to our problem setting, PartNeRF and DBW use images as input to construct structural 3D representations. To ensure fair comparisons, we discard the pre-defined ground plane in DBW, as the test scenario consists exclusively of foreground objects. For PartNeRF, we perform instance-specific training and set the same number of parts as in our method, while retaining all other default parameter settings. Additionally, we compare with SOTA Gaussian Splatting reconstruction method 2DGS [21], and surface reconstruction approach Neuralangelo [30], to provide an intuitive evaluation of the reconstruction quality achieved by our method.
6.3 Evaluation Metrics
We evaluate from 3D reconstruction quality, view synthesis, shape parsimony, and reconstruction time.
  • The 3D reconstruction quality is measured by the official Chamfer distance (CD) evaluation [23] between the recovered geometry and GT, reflecting the accuracy of 3D reconstructions.
  • View synthesis uses three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [57], and Learned Perceptual Image Patch Similarity (LPIPS) [69].
  • Shape parsimony is quantified by the number of parts, while reconstruction time refers to the running time measured on the same device. While shape parsimony does not directly indicate the quality of the decomposition, a smaller number of decomposed components, given consistent reconstruction quality, suggests a more concise and reasonable decomposition. This is because fitting different parts of an object with fewer blocks is inherently more challenging, whereas utilizing more blocks simplifies the fitting process.
Figure 10: Qualitative comparisons on DTU [23]. We compare our approach with state-of-the-art baselines on the DTU dataset with the background removed. The last two columns show our block-level and point-level reconstructions, respectively. Our method is the only one that provides reasonable 3D part decomposition while capturing detailed geometry.
Figure 11: Qualitative comparisons on ShapeNet [23]. We compared our approach with state-of-the-art baselines across four categories. The last two columns display our block-level and point-level reconstructions, respectively. Our method uniquely provides reasonable 3D part decomposition while simultaneously capturing detailed geometry.
Figure 12: Qualitative results on BlendedMVS [63] and self-capatured data. We demonstrate the RGB renderings and decomposed parts from novel views across a variety of objects. The first five examples are from the BlendedMVS dataset, and the remaining examples are from our own captured scenes.
7 Additional Comparative Experiments
Very recently, DPA-Net [68] and GaussianBlock [24] have achieved advanced 3D shape abstraction. DPA-Net enables part-aware reconstruction in a feedforward manner, requiring approximately 3 days of GPU training on a pre-collected dataset and 2 hours per object for inference. GaussianBlock utilizes SAM [27] to guide superquadric splitting and fusion for 3D decomposition, with a processing time of 6 hours per object. We compare our method on DTU and BlendedMVS scenes, with results presented in Fig. 13. The proposed approach achieves comparable performance while demonstrating superior efficiency. It is important to clarify the supervision requirements: DPA-Net is a supervised method that relies on SAM-generated segmentation masks for training, whereas GaussianBlock, while not strictly 3D supervised, depends on high-quality pre-trained datasets that typically require days of training. Our method distinguishes itself by being completely self-supervised, requiring neither 3D supervision nor pre-trained components, yet achieving significantly better computational efficiency.
Figure 13: Qualitative comparisons to DPA-Net and GaussianBlock. The first two examples are from the DTU dataset, and the last two examples are from the BlendedMVS dataset.
8 Implementation Details
8.1 Training Configurations
To ensure uniform distributions of 2D Gaussians across surfaces of the mesh, random barycentric coordinates are generated directly within each triangular face. Specifically, barycentric weights are computed as $u=\sqrt{rand}$ and $v=rand$. These weights are transformed to obtain $\alpha=[1-u,u(1-v),uv]$, $rand\in(0,1)$. The sampled position on triangular face is calculated as $o=\alpha_{0}v0+\alpha_{1}v1+\alpha_{2}v2$, where $v_{o},v_{1},v_{2}$ are vertices of the triangle. This method compensates for the non-uniformity caused by the triangle’s geometry. Unlike naive random sampling which results in uneven distributions, it ensures that the sampled points are evenly distributed across mesh surfaces.
The same hyperparameters are used for all experiments. We set the initial number of primitives $M$ to 8. In the hybrid representation, each superquadric mesh is a level-2 icosphere (320 triangular face). Each triangular face contains 100 Gaussians, with a scaling parameter $c$ of 0.1. The number of sampled points in each ray is 2048.
During refinement, we employ regularization in 2DGS to achieve better geometric reconstruction, including depth distortion maps, depth maps, and normal maps. Additionally, we introduce a mask cross-entropy to filter out extra noise Gaussians. To extract the meshes from 2D Gaussians, we use truncated signed distance fusion (TSDF) to fuse rendered depth maps, utilizing Open3D [71].
8.2 Rendering
With Gaussians attached to the surface of each block, we achieve view-dependent rendering through tile-based rasterization as Gaussian Splatting [25, 21]. Given a view, Gaussians on the superquadric surface are projected onto the image space, forming an RGB image. Initially, the screen space determines the bounding box for each Gaussian. Subsequently, these Gaussian ellipses are sorted according to their depths of center to the image plane. Finally, volumetric alpha compositing [41] is utilized to integrate the alpha-weighted RGB values for each pixel.
To formulate the process, considering a viewing transformation ${W}$, the covariance matrix ${\Sigma}_{i}^{{}^{\prime}}$ of $i$-th Gaussian in the camera coordinate system is calculated by:
${\Sigma}_{i}^{{}^{\prime}}={J}{W}{\Sigma}{W}^{T}{J}^{T},$ (14)
where ${J}$ is the Jacobian of the affine approximation of the projective transformation, and ${\Sigma}$ is the covariance matrix of the Gaussian ellipse. Note that the last row and column of ${\Sigma}$ are omitted since we adopt 2D Gaussians. Following alpha compositing, we first calculate an alpha value for each Gaussian ellipse:
$\alpha_{i}=\tau_{i}\exp(-\frac{1}{2}(\mathbf{x}-\mathbf{u}_{i})^{T}{\Sigma_{i}^{{}^{\prime}}}_{-1}(\mathbf{x}-\mathbf{u}_{i}))).$ (15)
Here, $\mathbf{u}_{i}$ is the center coordinate of the projected Gaussian ellipse, and $\tau_{i}$ is the opacity of the block where the $i$-th Gaussian is located. The calculated alphas are sorted according to their depths from the image plane. Meanwhile, we can acquire the color value $c_{i}$ from the spherical harmonics of $\mathcal{N}$ ordered points, thus obtaining the rendering RGB value:
$C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (16)
This process is differentiable and can optimize the hybrid representation through gradient descent.